# Bayesian inversion with Student’s $t$ priors based on Gaussian scale mixtures

### v1

@@ 3.3 Posterior under the Student’s $t$ prior
This one-dimensional density can be sampled within the Gibbs iterations using classical algorithms such as rejection sampling or random walk Metropolis (RWM). In our study, after comparing both methods, we choose to use RWM. For this RWM-within-Gibbs sampling step we can also perform a few extra within-Gibbs iterations (i.e., a within-Gibbs burn-in $\bar{n}_{b}$ and thinning $\bar{n}_{t}$) in order to accelerate the convergence of the Gibbs chain and to reduce correlation in the resulting posterior samples (see, e.g., (Gamerman and Lopes, 2006, p.213)). To improve the convergence, we use vanishing adaptation during the within-Gibbs burn-in Andrieu and Thoms (2008). This type of adaptation ensures that the proposal depends on recent sates in the chain less and less strongly, i.e., the adaptation gradually disappears from the proposal.
4 Computation procedure for posterior sampling
We describe two MCMC algorithms to sample from the posterior distribution under the Student’s $t$ prior. The first algorithm, the Gibbs sampler, exploits the GSM representation of the Student’s $t$ distribution. The other algorithm, NUTS, uses the traditional Student’s $t$ formulation; we consider NUTS to validate the numerical results obtained using Gibbs sampling.
We remark that, from a practical perspective, in MCMC we generate $n_{s}$ states of the chain that are only selected after discarding $n_{b}$ samples during the warm up phase of the algorithm (burn-in period), and the sample chain values are only stored every $n_{t}$ iterations since for large enough $n_{t}$ the samples are virtually independent (lagging steps). As a result, we need to draw $n=n_{b}+n_{s}n_{t}$ Markov chain steps to obtain $n_{s}$ quasi-independent samples from the posterior. Furthermore, the quality of a chain produced by a sampler can be quantitatively measured using the effective sample size (ESS). This metric reflects the amount of uncertainty within the chain due to autocorrelation. Therefore, if $\rho^{(i)}$ denotes the autocorrelation at the $i$-th lag, the ESS can be calculated as follows
where $\tau_{\mathrm{int}}$ stands for the integrated autocorrelation time (IACT). The definition of the ESS intuitively means that the larger chain autocorrelation is, the smaller effective sample size is. Hence, while sampling one should target values of $n_{\text{eff}}$ as close as possible to the sample size $n_{s}$.
@@ 4.1 Gibbs sampler
Gibbs sampling operates as an iterative algorithm that systematically selects a single variable from a model and resamples it based on its conditional distribution, given the other variables. The sampling process is based on the fact that, under mild conditions, the set of full conditional distributions determines the joint distribution Besag (1974). Moreover, convergence conditions of the Gibbs sampler are discussed in Roberts and Smith (1994); Tierney (1994). By deriving the full conditional densities of all parameters, as shown in Section 3.3, we can use the Gibbs sampler to generate a Markov chain with stationary distribution equivalent to the posterior distribution eq. 16.
In Gibbs sampling, the process of choosing which parameter to sample (i.e., the variable index), often referred to as the scan order, plays a crucial role Owen (2023). Two predominant scan orders are employed: random scan and systematic scan. In random scan, variables are chosen uniformly and independently at random during each iteration. In systematic scan, a predetermined permutation is chosen, with variables consistently selected in that specified order throughout the sampling process. Without loss of generality we follow a random scan strategy; this is mainly due to the fact that this Gibbs sampler version is reversible with respect to the posterior. The resulting Gibbs sampler applied to posterior eq. 16 is summarized (with a slight abuse of notation) in Algorithm 1.
Data: Gibbs parameters: number of samples $n_{s}$, burn-in $n_{b}$, thinning $n_{t}$. RWM-within-Gibbs parameters: warm-up $\bar{n}_{b}$ and thinning $\bar{n}_{t}$.                               Posterior conditional densities section 3.3.
1 Initialize $\bm{x}$, $\sigma_{\mathrm{obs}}^{2}$, $\tau^{2}$, $\bm{w}^{2}$, $\nu$ $n\leftarrow n_{s}n_{t}+n_{b}$ $j\leftarrow 1$ for $i\leftarrow 1$ to $n$ do
2       $k\leftarrow\texttt{randint}([1,5])$   # generate the variable index
@@
5                   compute posterior mean $\widetilde{\bm{\mu}}(\tau^{2},\bm{w}^{2})$ and precision matrix $\widetilde{\bm{\Lambda}}\;(\tau^{2},\bm{w}^{2})$sample $\bm{x}\sim\mathcal{N}(\widetilde{\bm{\mu}},\widetilde{\bm{\Lambda}}^{-1})$
6             case $2$ do sample $\sigma_{\mathrm{obs}}^{2}\sim p(\cdot\;|\;\bm{x})$ in closed form according to eq. 20d  case $3$ do sample $\tau^{2}\sim p(\cdot\;|\;\bm{x},\bm{w}^{2})$ in closed form according to eq. 21d  case $4$ do sample $\bm{w}^{2}\sim p(\cdot\;|\;\bm{x},\tau^{2},\nu)$ in closed form according to eq. 22d  case $5$ do sample $\nu\sim p(\cdot\;|\;\bm{w}^{2})$ in eq. 23 using RWM with warm-up $\bar{n}_{b}$ and thinning $\bar{n}_{t}$
@@
9             $\bm{x}^{(j)}\leftarrow\bm{x}$, $(\sigma_{\mathrm{obs}}^{2})^{(j)}\leftarrow\sigma_{\mathrm{obs}}^{2}$, $(\tau^{2})^{(j)}\leftarrow\tau^{2}$, $(\bm{w}^{2})^{(j)}\leftarrow\bm{w}^{2}$, $\nu^{(j)}\leftarrow\nu$ $j\leftarrow j+1$
@@
4.2 NUTS sampler
We introduced the Bayesian inverse problem of estimating the posterior eq. 16 under the MRF prior based on the GSM representation of the Student’s $t$ distribution. The GSM formulation lends itself well to the application of the Gibbs sampler, due to the availability of conditional densities for each parameter.
Alternatively, we consider imposing Student’s $t$ distribution on the vector of increments $\bm{u}=\bm{L}\bm{x}$ without the GSM representation. In this case, the posterior eq. 6b can be modelled hierarchically as follows
@@ 4.2 NUTS sampler
$(\bm{L}\bm{x}\;|\;\tau^{2},\nu)\sim\mathcal{T}_{\nu}\left(\mu,\tau^{2}\right),$ (25b)
$\tau^{2}\sim\text{IG}(a,b),\quad\nu\sim p(\nu).$ (25c)
To sample from the posterior eq. 25a, we employ NUTS Hoffman and Gelman (2014). This method requires the logarithm of posterior and also its gradient. The logarithm of posterior eq. 25a is defined as follows
$\log p(\bm{x},\tau^{2},\nu\;|\;\bm{y})=\log p(\bm{y}\;|\;\bm{x})+\log p(\bm{L}\bm{x}\;|\;\tau^{2},\nu)+\log p(\tau^{2})+\log p(\nu).$ (26)
Since our formulation relies on linear forward models and the likelihood is Gaussian, the log-likelihood and its derivative with respect to the target parameter $\bm{x}$ are well-known (see, e.g., (Petersen and Pedersen, 2012, p.11)). However, as the prior distribution is more involved, we now write its natural logarithm and associated derivatives. The log-prior on $\bm{x}$ is
$\begin{split}\log p(\bm{L}\bm{x}\;|\;\tau^{2},\nu)\propto&\,d\left[\log\Gamma\left(\frac{\nu+1}{2}\right)-\log\Gamma\left(\frac{\nu}{2}\right)-\frac{1}{2}\log(\pi\nu)-\frac{1}{2}\log(\tau^{2})\right]-\\ &\left(\frac{\nu+1}{2}\right)\sum_{i=1}^{d}\log\left(1+\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{\nu\tau^{2}}\right),\end{split}$ (27)
@@
We now follow the same procedure for the hyperparameters, assuming a gamma prior $\text{Ga}(\nu;\,a,b)$ on the degrees of freedom parameter. The logarithms and derivatives of the hyperprior for $\tau^{2}$ and prior for $\nu$ are defined as
$\log p(\tau^{2})$ $\propto(-a-1)\log(\tau^{2})-\frac{b}{\tau^{2}},\quad$ $\frac{\partial}{\partial\tau^{2}}\log p(\tau^{2})$ $\propto-\frac{a+1}{\tau^{2}}+\frac{b}{(\tau^{2})^{2}},$ (29a)
$\log p(\nu)$ $\propto(a-1)\log(\nu)-b\,\nu,\quad$ $\frac{\partial}{\partial\nu}\log p(\nu)$ $\propto\frac{a-1}{\nu}-b.$ (29b)
@@ 5 Numerical experiments
In the following, we illustrate the use of the proposed Student’s $t$ priors for Bayesian inversion. We consider linear inverse problems arising in signal processing and imaging science. The first example is one-dimensional deconvolution problem that allows us to test multiple parameter settings in our computational framework. The second example is the image deblurring problem.
In our test problems, the point estimates $\bar{\bm{x}}$ (posterior mean) of the target parameter are evaluated using the relative reconstruction error defined as
$\text{relerr}:=\frac{\left\|\bar{\bm{x}}-\bm{x}^{\mathrm{true}}\right\|_{2}}{\left\|\bm{x}^{\mathrm{true}}\right\|_{2}},$ (30)
where $\bm{x}^{\mathrm{true}}$ denotes the underlying true solution.
@@ 5.1 One-dimensional deconvolution
A finite-dimensional representation of eq. 31 is employed in practice. After discretizing the signal domain using $d=128$ points, the convolution model can be expressed as a system of linear algebraic equations $\bm{y}=\bm{A}\bm{x}$. We consider synthetic observation data $\bm{y}$ with $m=d-2$ equally-spaced elements (excluding two boundary points).
We compare two different signals, a piecewise constant (or sharp) signal and smooth signal, to illustrate the flexibility of the proposed priors and their performance in recovering discontinuous signals as well as smooth. Figure 2 shows the two underlying true signals together with the noisy observational data. The data were obtained according to model eq. 31 with parameter $\sigma=4$ for the piecewise constant signal and $\sigma=8$ for smooth. The data were corrupted by adding Gaussian noise with standard deviation $\sigma_{\mathrm{obs}}^{\mathrm{true}}=8.654\times 10^{-3}$ for the piecewise constant signal and $\sigma_{\mathrm{obs}}^{\mathrm{true}}=4.368\times 10^{-2}$ for the smooth signal.
Figure 2: One-dimensional deconvolution example. Left: sharp signal; right: smooth signal. Noisy convolved data (dots).
Firstly, we perform a prior study for the degrees of freedom parameter $\nu$ to select the best suitable option for our setting. Next, for the selected prior, we compare Gibbs and NUTS samplers and their efficiency in exploring the posterior distribution. Finally, we compare solutions obtained with the Student’s $t$ prior against the results obtained using another popular prior — the Laplace prior.
@@
We explore the resulting posterior distribution using the Gibbs sampler with the number of samples $n_{s}=2\times 10^{4}$, the global burn-in (or warm-up) period $n_{b}=2\times 10^{3}$, and lag (or thinning number) $n_{t}=20$. To sample the degrees of freedom parameter $\nu$ within Gibbs, we use RWM sampler that produces $1$ sample with burn-in period $\bar{n}_{b}=100$ and thinning $\bar{n}_{b}=10$. We choose values for $\bar{n}_{b}$ and $\bar{n}_{t}$ based on numerical studies. Since conditional of $\nu$ eq. 23 is computationally inexpensive to evaluate we can afford to perform extra within-Gibbs iterations (burn-in and thinning) to improve the resulting sub-chain of $\nu$.
@@
In Table 1, we report posterior statistics (sample mean, standard deviation and effective sample size) for the scalar parameters: degrees of freedom $\nu$, global scale parameter $\tau$ and the noise standard deviation $\sigma_{\mathrm{obs}}$.
Based on the analysis, we choose $\text{Ga}_{\nu>1}(\nu;\,2,0.1)$ as a prior distribution for the degrees of freedom parameter. This choice results in good reconstructions for both discontinuous and smooth cases and, in addition, allows avoiding possible pathological behavior via restricting values of $\nu$. We remark that the values of the proposal scaling in the RWM-within-Gibbs after adaptation are around $0.555$ and $1.067$, for the sharp and smooth signals, respectively.
@@
Sampler comparison: Gibbs vs. NUTS. Using the prior distribution $\text{Ga}_{\nu>1}(\nu;\,2,0.1)$ on the degrees of freedom parameter $\nu$, we compare posterior statistics of $\nu$ obtained using the Gibbs sampler with those obtained via the NUTS sampler. In NUTS, we follow the standard Student’s $t$ formulation $\mathcal{T}_{\nu}\left(\mu,\tau^{2}\right)$ with hyperpriors $\tau^{2}\sim\text{IG}(1,10^{-4})$ and $\nu\sim\text{Ga}_{\nu>1}(2,0.1)$ as described in Section 4.2. We run NUTS sampler to obtain $n_{s}=2\times 10^{4}$ states and apply additionally burn-in of $n_{b}=2\times 10^{3}$.
Signal reconstructions are of the same good quality when either Gibbs or NUTS is used. In Figure 6, we compare the posterior statistics for the parameter $\nu$ obtained using both samplers for the sharp signal. Since the posterior chains are comparatively good, we can conclude that the Gibbs sampler is exploring the target posterior distribution. We omit the analogous chain plot for the smooth signal scenario in the sake of brevity. However, we note that inference of the $\nu$ parameter is more challenging in the smooth signal case, as the Gibbs chain has slightly more correlation than the one computed by NUTS. Nonetheless, we remark here that the main advantage of the Gibbs sampler over NUTS is its independence from gradient computations.
@@
Prior comparison: Student’s $t$ prior vs. Laplace prior. We also compare posterior obtained with the Students’s $t$ prior with the posterior obtained using the Laplace prior Bardsley (2012). In one-dimensional case, Laplace prior is equivalent to the total variation prior and is defined as
$p(\bm{x})=\left(\frac{\delta}{2}\right)^{d}\exp\left(-\delta\|\bm{D}\bm{x}\|_{1}\right),$ (32)
where the difference matrix $\bm{D}$ is defined in eq. 4, parameter $\delta\in\mathbb{R}_{>0}$ is the inverse scale, and $\left\|\cdot\right\|_{1}$ denotes the $\ell_{1}$-norm. We sample both cases using NUTS ($n_{s}=2\times 10^{4}$ states with additional burn-in $n_{b}=2\times 10^{3}$).
In Figure 7, we show posterior statistics of parameter $\bm{x}$. In the smooth signal test, both prior behave comparatively well resulting in small relative error and similar uncertainty. However, in the piecewise constant case, our Student’s $t$ prior improves the reconstruction error and reduces the posterior uncertainty.
Figure 7: Comparison of posterior statistics for the target parameter $\bm{x}$ using NUTS. Left: Student’s $t$ prior. Right: Laplace prior. Results are shown for the sharp signal (top row) and smooth signal (bottom row). True solution (black solid line), posterior median (red dashed line), posterior mean (blue dotted line), and $95\%$ highest density interval (shaded area).

### v2

@@ 3.3 Posterior under the Student’s $t$ prior
This one-dimensional density can be sampled within the Gibbs iterations using classical algorithms such as rejection sampling or random walk Metropolis (RWM). In our study, after comparing both methods, we choose to use RWM. For this RWM-within-Gibbs sampling step, we can also perform a few extra within-Gibbs iterations (i.e., a within-Gibbs burn-in $\bar{n}_{b}$) in order to accelerate the convergence of the Gibbs chain and to reduce the correlation in the resulting posterior samples (see, e.g., [49, p.213]). To improve the sampling process, we use vanishing adaptation during the within-Gibbs burn-in [50]. This type of adaptation ensures that the proposal is less and less dependent on the recent states in the chain, i.e., the adaptation gradually disappears from the proposal.
4 Computational procedure for posterior sampling
We describe two MCMC algorithms to sample from the posterior distribution under the Student’s $t$ prior. The first algorithm, the Gibbs sampler, exploits the GSM representation of Student’s $t$ distribution. The other algorithm, NUTS, uses the traditional Student’s $t$ formulation; we consider NUTS to validate the numerical results obtained using Gibbs sampling.
We remark that, from a practical perspective, we generate $n_{s}$ samples from the Markov chain, which are selected after discarding $n_{b}$ samples during the warm-up phase of the algorithm (so-called burn-in period). Moreover, to reduce storage requirements and improve the sample correlation, the sample chain values can be stored every $n_{t}$ iterations (lagging steps). Therefore, we need to draw $n=n_{b}+n_{s}n_{t}$ Markov chain steps to obtain $n_{s}$ samples from the posterior.
@@ 4.1 Gibbs sampler
Gibbs sampling operates as an iterative algorithm that systematically selects a single variable and resamples it based on its conditional distribution, given the other variables. The sampling process is based on the fact that, under mild conditions, the set of full conditional distributions determines the joint distribution [51]. Moreover, convergence conditions of the Gibbs sampler are discussed in [52, 53]. By deriving the full conditional densities of all parameters, as shown in Section 3.3, we can use the Gibbs sampler to generate a Markov chain with a stationary distribution equivalent to the posterior distribution eq. 16.
In Gibbs sampling, the process of choosing which parameter to sample (i.e., the variable index), often referred to as the scan order, plays a crucial role [54]. Two predominant scan orders are employed: random scan and systematic scan. In random scan, variables are chosen uniformly and independently at random during each iteration. In systematic scan, a predetermined permutation is chosen, with variables consistently selected in that specified order throughout the sampling process. Without loss of generality, we follow a random scan strategy; this is mainly due to the fact that this Gibbs sampler version produces a Markov chain that is reversible with respect to the posterior. The resulting Gibbs sampler applied to posterior eq. 16 is summarized (with a slight abuse of notation) in Algorithm 1.
Data: Gibbs parameters: number of samples $n_{s}$, burn-in $n_{b}$, thinning $n_{t}$. RWM-within-Gibbs parameters: warm-up $\bar{n}_{b}$.
            Posterior conditional densities eq. 17.
1 Initialize $\bm{x}$, $\sigma_{\mathrm{obs}}^{2}$, $\tau^{2}$, $\bm{w}^{2}$, $\nu$
2 $n\leftarrow n_{s}n_{t}+n_{b}$
3 $j\leftarrow 1$
4 for $i\leftarrow 1$ to $n$ do
5       $k\leftarrow\texttt{randint}([1,5])$
6        # generate the variable index
@@ 5       $k\leftarrow\texttt{randint}([1,5])$
9                   compute posterior mean $\widetilde{\bm{\mu}}(\tau^{2},\bm{w}^{2})$ and precision matrix $\widetilde{\bm{\Lambda}}\;(\tau^{2},\bm{w}^{2})$
10                  sample $\bm{x}\sim\mathcal{N}(\widetilde{\bm{\mu}},\widetilde{\bm{\Lambda}}^{-1})$
11
12             case $2$ do sample $\sigma_{\mathrm{obs}}^{2}\sim p(\cdot\;|\;\bm{x})$ in a closed form according to eq. 20
13              case $3$ do sample $\tau^{2}\sim p(\cdot\;|\;\bm{x},\bm{w}^{2})$ in a closed form according to eq. 21
14              case $4$ do sample $\bm{w}^{2}\sim p(\cdot\;|\;\bm{x},\tau^{2},\nu)$ in a closed form according to eq. 22
15              case $5$ do sample $\nu\sim p(\cdot\;|\;\bm{w}^{2})$ in eq. 23 using RWM with warm-up $\bar{n}_{b}$
16
@@
19             $\bm{x}^{(j)}\leftarrow\bm{x}$, $(\sigma_{\mathrm{obs}}^{2})^{(j)}\leftarrow\sigma_{\mathrm{obs}}^{2}$, $(\tau^{2})^{(j)}\leftarrow\tau^{2}$, $(\bm{w}^{2})^{(j)}\leftarrow\bm{w}^{2}$, $\nu^{(j)}\leftarrow\nu$
20            $j\leftarrow j+1$
21
@@ 20            $j\leftarrow j+1$
4.2 NUTS algorithm
We introduced the Bayesian inverse problem of estimating the posterior eq. 16 under the MRF prior, based on the GSM representation of Student’s $t$ distribution. The GSM formulation lends itself well to the application of the Gibbs sampler, due to the availability of conditional densities for each parameter.
Alternatively, we consider imposing Student’s $t$ distribution on the vector of increments $\bm{u}=\bm{L}\bm{x}$ without the GSM representation. In this case, the posterior eq. 6b can be modeled hierarchically as follows
@@ 4.2 NUTS algorithm
$(\bm{L}\bm{x}\;|\;\tau^{2},\nu)\sim\mathcal{T}_{\nu}\left(\mu,\tau^{2}\right),\quad\tau^{2}\sim\text{IG}(a,b),\quad\nu\sim p(\nu);$ (24b)
to sample from this posterior, we can employ Hamiltonian Monte Carlo based on its No-U-Turn Sampler adaptation [34]. This method requires the logarithm of the posterior and also its gradient.
Our formulation relies on linear forward models and Gaussian likelihoods, hence, the log-likelihood and its derivative with respect to the target parameter $[\bm{x},\tau^{2},\nu]$ are well-known (see, e.g., [55, p.11]). However, as the prior distribution is more involved, we now write its logarithm and associated derivatives. The log-prior on $\bm{x}$ is
$\begin{split}\log p(\bm{L}\bm{x}\;|\;\tau^{2},\nu)\propto&\,d\left[\log\Gamma\left(\frac{\nu+1}{2}\right)-\log\Gamma\left(\frac{\nu}{2}\right)-\frac{1}{2}\log(\pi\nu)-\frac{1}{2}\log(\tau^{2})\right]-\\ &\left(\frac{\nu+1}{2}\right)\sum_{i=1}^{d}\log\left(1+\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{\nu\tau^{2}}\right),\end{split}$
@@
Finally, assuming a gamma prior $\text{Ga}(\nu;\,a,b)$ on the degrees of freedom parameter, we can follow the same procedure for the hyperparameters. The logarithms and derivatives of the hyperprior for $\tau^{2}$ and prior for $\nu$ are defined as
$\log p(\tau^{2})$ $\propto(-a-1)\log(\tau^{2})-\frac{b}{\tau^{2}},\quad$ $\frac{\partial}{\partial\tau^{2}}\log p(\tau^{2})$ $\propto-\frac{a+1}{\tau^{2}}+\frac{b}{(\tau^{2})^{2}},$
$\log p(\nu)$ $\propto(a-1)\log(\nu)-b\,\nu,\quad$ $\frac{\partial}{\partial\nu}\log p(\nu)$ $\propto\frac{a-1}{\nu}-b.$
@@ 5 Numerical experiments
In the following, we illustrate the use of the proposed Student’s $t$ priors for Bayesian inversion. We consider linear inverse problems arising in signal processing and imaging science. The first example is a one-dimensional deconvolution problem that allows us to test multiple parameter settings in our computational framework. The second example is image deblurring.
In our test problems, the point estimates $\bar{\bm{x}}$ (such as the posterior mean) of the target parameter are evaluated using the relative reconstruction error defined as
$\varepsilon_{\mathrm{rel}}:=\frac{\left\|\bar{\bm{x}}-\bm{x}^{\mathrm{true}}\right\|_{2}}{\left\|\bm{x}^{\mathrm{true}}\right\|_{2}},$ (27)
where $\bm{x}^{\mathrm{true}}$ denotes the underlying true solution. Furthermore, the quality of the MCMC samples is measured by the effective sample size (ESS). This metric reflects the amount of uncertainty within the chain due to autocorrelation. Therefore, if $\rho^{(i)}$ denotes the autocorrelation at the $i$-th lag, the ESS can be calculated as follows
where $\tau_{\mathrm{int}}$ stands for the integrated autocorrelation time (IACT). The definition of the ESS intuitively means that the larger the chain autocorrelation is, the smaller ESS is. Hence, while sampling one should target values of $n_{\text{eff}}$ as close as possible to the sample size $n_{s}$.
@@ 5.1 One-dimensional deconvolution
A finite-dimensional representation of Equation eq. 29 is employed in practice. After discretizing the signal domain using $d=130$ points, the convolution model can be expressed as a system of linear algebraic equations $\bm{y}=\bm{A}\bm{x}$. We consider synthetic observation data $\bm{y}$ with $m=d-2$ equally-spaced elements (excluding two boundary points).
We compare two different signals, a piecewise constant (or sharp) signal and a smooth signal, to illustrate the flexibility of the proposed priors and their performance in recovering discontinuous signals as well as smooth. Figure 2 shows the two underlying true signals together with the noisy observational data. The data were obtained according to model eq. 29 with parameter $\sigma=4$ for the piecewise constant signal and $\sigma=8$ for smooth. The data were corrupted by adding Gaussian noise with standard deviation $\sigma_{\mathrm{obs}}^{\mathrm{true}}=8.654\times 10^{-3}$ for the piecewise constant signal and $\sigma_{\mathrm{obs}}^{\mathrm{true}}=4.368\times 10^{-2}$ for the smooth signal.
Figure 2: One-dimensional deconvolution example. Left: sharp signal; right: smooth signal.
First, we perform a prior study for the degrees of freedom parameter $\nu$ to select the most suitable option for our setting. Next, for the selected prior, we compare Gibbs and NUTS samplers and their efficiency in exploring the posterior distribution. Finally, we compare solutions obtained with the Student’s $t$ prior against the results obtained using two other popular priors — the Laplace prior and the Cauchy difference prior.
@@
We explore the resulting posterior distribution using the Gibbs sampler with the number of samples $n_{s}=2\times 10^{4}$, the global burn-in (or warm-up) period $n_{b}=2\times 10^{3}$, and lag (or thinning number) $n_{t}=20$. To sample the degrees of freedom parameter $\nu$ within Gibbs, we use RWM sampler that produces one sample with burn-in period $\bar{n}_{b}=100$; we choose this value based on numerical studies. Since the conditional of $\nu$ in eq. 23 is computationally inexpensive to evaluate we can afford to perform extra within-Gibbs iterations (burn-in) to improve the resulting sub-chain of $\nu$.
@@
In Table 2, we report posterior statistics (sample mean, standard deviation, and effective sample size) for the scalar parameters: degrees of freedom $\nu$, global scale parameter $\tau$, and the noise standard deviation $\sigma_{\mathrm{obs}}$. Based on this analysis, we choose $\text{Ga}_{\nu>1}(\nu;\,2,0.1)$ as a prior distribution for the degrees of freedom parameter. This choice results in good reconstructions for both discontinuous and smooth cases and, in addition, allows for avoiding possible pathological behavior via restricting values of $\nu$. We remark that the values of the proposal scaling in the RWM-within-Gibbs after adaptation are around $0.555$ and $1.067$, for the sharp and smooth signals, respectively.
@@
Sampler comparison: Gibbs vs. NUTS. Using the prior distribution $\text{Ga}_{\nu>1}(\nu;\,2,0.1)$ on the degrees of freedom parameter $\nu$, we compare posterior statistics of $\nu$ obtained via the Gibbs sampler with those obtained via NUTS. In NUTS, we follow the standard formulation of the $t$ distribution $\mathcal{T}_{\nu}\left(\mu,\tau^{2}\right)$ with hyperpriors $\tau^{2}\sim\text{IG}(1,10^{-4})$ and $\nu\sim\text{Ga}_{\nu>1}(2,0.1)$, as described in Section 4.2. We run NUTS to obtain $n_{s}=2\times 10^{4}$ states and apply additionally burn-in of $n_{b}=2\times 10^{3}$.
Signal reconstructions are of the same good quality when either Gibbs or NUTS is used. In Figure 6, we compare the posterior statistics for the parameter $\nu$ obtained using both samplers for the sharp signal. Since the posterior chains are comparatively good, we can conclude that the Gibbs sampler is appropriately exploring the target posterior distribution. We omit the analogous chain plot for the smooth signal scenario for the sake of brevity. However, we note that inference of the $\nu$ parameter is more challenging in the smooth signal case, as the Gibbs chain has slightly more correlation than the one computed by NUTS. Nonetheless, we remark here that the main advantage of the Gibbs sampler over NUTS is its independence from gradient computations.
@@
Prior comparison: Student’s $t$ prior vs. Laplace prior and Cauchy prior. We also compare the posterior obtained using the Student’s $t$ prior with the posteriors obtained using the Laplace prior [23] and the Cauchy difference prior [56]. In the one-dimensional case, the Laplace and Cauchy difference priors are defined as
$p(\bm{x})=\begin{cases}\left(\dfrac{1}{2\tau}\right)^{d}\exp\left(-\dfrac{\left\|\bm{D}\bm{x}\right\|_{1}}{\tau}\right)&\text{Laplace}\\ \left(\dfrac{1}{\pi}\right)^{d}\displaystyle\prod\limits_{i=1}^{d}\dfrac{\tau}{\left([\bm{D}\bm{x}]_{i}^{2}+\tau^{2}\right)}&\text{Cauchy},\end{cases}$ (30)
where the difference matrix $\bm{D}$ is specified in eq. 4, the parameter $\tau\in\mathbb{R}_{>0}$ is the scale, and $\left\|\cdot\right\|_{1}$ denotes the $\ell_{1}$-norm. Recall that the Cauchy difference prior is equivalent to the Student’s $t$ MRF prior with fixed $\nu=1$. In this study, we sample both cases using NUTS ($n_{s}=2\times 10^{4}$ states with additional burn-in $n_{b}=2\times 10^{3}$).
In Figure 7, we show posterior statistics of parameter $\bm{x}$ for Student’s $t$, Laplace, and Cauchy priors. In the piecewise constant case, the Student’s $t$ prior results in a similar reconstruction quality as compared to the Cauchy prior; whereas the $t$ prior clearly reduces the reconstruction error and the posterior uncertainty as compared to the Laplace prior. In the smooth signal case, the Student’s $t$ prior outperforms the Cauchy prior due to its flexibility, which allows the reconstruction of smooth features as well as sharp ones.
Figure 7: Comparison of posterior statistics for the target parameter $\bm{x}$ using NUTS. Left: Student’s $t$ prior, middle: Laplace prior, right: Cauchy prior. Results are shown for the sharp signal (top row) and smooth signal (bottom row). True solution (black solid line), posterior median (red dashed line), posterior mean (blue dotted line), and $95\%$ highest density interval (shaded area).
Remark 5.1.
We also studied the influence of our Student’s $t$ MRF prior on the convergence of the sampling method under different discretization sizes in comparison to the Laplace prior. We used the sharp signal data set ($m=128$), considered sizes $d=\{64,128,256\}$, and estimated the posterior using NUTS ($n_{s}=2\times 10^{4}$ and $n_{b}=2\times 10^{3}$) for both priors. In each case, we computed five independent chains to check convergence using the estimated potential scale reduction factors (so-called $\widehat{R}$ diagnostic test) [57]. For each discretization size and all the sample chains, we obtained $\widehat{R}$ values close to one for both priors. Thus, we assume that all our sample chains have converged. While the solutions may show some variation with different discretization sizes, this does not affect the convergence of the sampler.